# A Survey of Mamba

### v2

@@ Abstract.
Deep learning, as a vital technique, has sparked a notable revolution in artificial intelligence (AI), resulting in a great change in human lifestyles. As the most representative architecture, Transformers have empowered numerous advanced models, especially the large language models (LLMs) that comprise billions of parameters, becoming a cornerstone in deep learning. Despite the impressive achievements, Transformers still face inherent limitations, particularly the time-consuming inference resulting from the quadratic computation complexity of attention calculation. Recently, a novel architecture named Mamba, drawing inspiration from classical state space models, has emerged as a promising alternative for building foundation models, delivering comparable modeling abilities to Transformers while preserving near-linear scalability concerning sequence length. This has sparked an increasing number of studies actively exploring Mamba’s potential to achieve impressive performance across diverse domains. Given such rapid evolution, there is a critical need for a systematic review that consolidates existing Mamba-empowered models, offering a comprehensive understanding of this emerging model architecture. In this survey, we therefore conduct an in-depth investigation of recent Mamba-associated studies, covering from three main aspects: the advancements of Mamba-based models, the techniques of adapting Mamba to diverse data, and the applications where Mamba can excel. Specifically, we first recall the foundational knowledge of various representative deep learning models and the details of Mamba-1&2 as preliminaries. Then, to showcase the significance of Mamba for AI, we comprehensively review the related studies focusing on Mamba models’ architecture design, data adaptability, and applications. Finally, we present an discussion of current limitations and explore various promising research directions to provide deeper insights for future investigations.
State Space Model, Mamba, Sequence Modeling, Foundation Models, Language Models
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2024; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Over the past two decades, deep learning, as the most prominent artificial intelligence (AI) technique, has brought about a revolution in various domains such as healthcare (Jones et al., 2024), autonomous systems (Guan et al., 2024; Fan et al., 2024a), recommender systems (Li et al., 2024a; Zhao et al., 2024b), and financial services (Prata et al., 2024; Zhang et al., 2024d). This period has witnessed the emergence of numerous deep neural networks (DNNs) that have significantly altered human lifestyles, offering immense convenience to individuals. One notable example is U-Net (Ronneberger et al., 2015; Si et al., 2024), a robust deep learning model within the field of vision, which is extensively employed in medical imaging for the examination of radiology scans like MRI and CT scans. Its application assists in the identification and diagnosis of diseases, showcasing its effectiveness in this critical healthcare domain (Williams et al., 2024; Lin et al., 2022). Moreover, Graph Neural Networks (GNNs) are employed in handling graph-structure data to support smart services, such as recommender systems that suggest personalized content, products, or services to users (Fan et al., 2020, 2019b; Wu et al., 2019). Furthermore, Recurrent Neural Networks (RNNs) are extensively adopted in machine translation due to their ability to capture the sequential and contextual information essential for accurate translations (Liu et al., 2014; Su et al., 2017), empowering individuals from diverse linguistic backgrounds to effectively communicate and comprehend each other’s ideas, opinions, and information.
@@
Among the various deep learning architecture, Transformers have recently stood out and established its dominance across a broad spectrum of applications (Dong et al., 2023; Vert, 2023). For instance, as the most representative large foundation models, large language models (LLMs) like ChatGPT are fundamentally built on the Transformer architecture (Achiam et al., 2023; Qu et al., 2024a; Zhao et al., 2024b). By scaling their model sizes to billions and training on a mix of diverse data sources, these Transformer-based models have demonstrated human-level intelligence with its impressive capabilities in language understanding, common sense reasoning, and in-content-learning (Zhang et al., 2023; Fan et al., 2024b). This remarkable success is bolstered by the attention mechanism (Vaswani et al., 2017), which enables the Transformer-based models to concentrate on relevant parts of the input sequence and facilitate better contextual understanding. However, the attention mechanism also introduces a significant computational overhead that increases quadratically with the input size (Lu et al., 2021; Zhu et al., 2021), which presents challenges in processing lengthy inputs. For example, the rapid growth in computational cost makes Transformers impractical or infeasible to process substantial sequences, thereby limiting their applicability in tasks like document-level machine translation (Maruf et al., 2021) or long document summarization (Koh et al., 2022).
@@
Recently, a promising architecture, structured state space sequence models (SSMs) (Gu et al., 2021a), have emerged to efficiently capture complex dependencies in sequential data, becoming a formidable competitor to Transformer. These models, inspired by classical state space models (Kalman, 1960), can be considered as a fusion of recurrent neural networks and convolutional neural networks. They can be computed efficiently using either recurrence or convolution operations, achieving linear or near-linear scaling with sequence length, thus significantly reducing the computational costs. More specifically, as one of the most successful SSM variants, Mamba achieves comparable modeling capabilities to Transformers while maintaining linear scalability with sequence length (Gu and Dao, 2023), propelling it into the realm of focal topics. Mamba first introduces a simple yet effective selection mechanism that enables the model to filter out irrelevant information while retaining necessary and relevant data indefinitely by parameterizing the SSM parameters based on the input. Then, Mamba proposes a hardware-aware algorithm to compute the model recurrently with a scan instead of convolution, achieving up to 3$\times$×faster computation on A100 GPUs. As shown in Figure 1, the powerful modeling capabilities for complex and lengthy sequential data, along with near-linear scalability, position Mamba as an emerging foundation model, poised to revolutionize various domains of research and applications, such as computer vision (Xu et al., 2024b; Zhu et al., 2024), natural language processing (Lieber et al., 2024; Zhao et al., 2024c), healthcare (Ruan and Xiang, 2024; Xing et al., 2024; Wang et al., 2024g), etc. For example, Zhu et al. (2024) propose Vim, which is 2.8$\times$×faster than DeiT (Touvron et al., 2021) and saves 86.8% GPU memory when extracting features for high-resolution images. Dao and Gu (2024) show the connections between SSMs and variants of attention and propose a new architecture that refines selective SSM, achieving 2-8$\times$ faster on language modeling.
@@ 2. Preliminary
Mamba is deeply intertwined with the recurrent framework of Recurrent Neural Networks (RNNs), the parallel computation and attention mechanism of Transformers, and the linear property of State Space Models (SSMs). Therefore, this section aims to present an overview of these three prominent architectures, serving as a prerequisite for comprehending Mamba thoroughly.
2.1. Recurrent Neural Networks
Recurrent Neural Networks (RNNs) excel in processing sequential data due to their capability to retain internal memory (Graves and Graves, 2012). Such networks have demonstrated remarkable effectiveness in a multitude of tasks that involve analyzing and predicting sequences, e.g., speech recognition, machine translation, natural language processing, and time-series analysis (Sutskever et al., 2011; Hermans and Schrauwen, 2013). In order to grasp the foundations of recurrent models, this section will offer a brief overview of the standard RNN formulation.
@@ 2.1. Recurrent Neural Networks
However, there are several limitations associated with RNNs. First, RNNs have a restricted capability to effectively extract long-range dynamics within input sequences. As information traverses through successive time steps, the repeated multiplication of weights in the network can lead to dilution or loss of information. Consequently, it becomes challenging for RNNs to retain and recall information from earlier time steps while making predictions. Second, RNNs process sequential data incrementally, restricting their computational efficiency since each time step relies on the preceding one. This makes parallel computations challenging for them. Furthermore, conventional RNNs lack built-in attention mechanisms, which allow the network to capture global information within input sequences. This absence of attention mechanisms hinders the network’s ability to selectively model the crucial segments of the data. o overcome these constraints, Transformers and State Space Models have emerged, each tackling these challenges from different perspectives. These two approaches will be further elaborated upon in the subsequent subsections.
@@ 2.2. Transformers
The Transformer (Vaswani et al., 2017) is a groundbreaking model in the realm of deep learning, revolutionizing various artificial intelligence applications. Its introduction marked a significant departure from traditional sequence-to-sequence models by employing a self-attention mechanism, facilitating the capture of global dependencies within model inputs. For instance, in natural language processing, this self-attention capability allows the model to comprehend relationships between various positions in a sequence. It achieves this by assigning weights to each position based on its significance relative to other positions. More specifically, a sequence of input vectors x is first transformed into three types of vectors: Query $Q$, Key $K$, and Value $V$ by utilizing linear transformations of the original input, defined by:
@@
Apart from performing a single attention function, multi-head attention is introduced to enhance the model’s ability to capture different types of relationships and provide multiple perspectives on the input sequence. In multi-head attention, an input sequence is processed in parallel by multiple sets of self-attention modules. Each head operates independently, performing the exact computations as in the standard self-attention mechanism. The attention weights from each head are then combined to create a weighted sum of the value vectors. This aggregation step allows the model to leverage information from multiple heads and capture diverse patterns and relationships present in the input sequence. Mathematically, the multi-head attention is computed as follows:
(5) $\begin{split}\text{MultiHead}&(Q,K,V)=(S_{1}\oplus S_{2}...\oplus S_{m})\cdot\mathbf{W}^{O},\\ \text{where}~{}~{}~{}~{}&S_{i}=\text{Softmax}(\frac{Q_{i}K_{i}^{T}}{\sqrt{d_{K}}})V_{i},~{}~{}i\in[1,m],\end{split}$
where $m$ is the number of attention heads. $\oplus$ is the concatenation operation, and $\mathbf{W}^{O}$ is the linear transformation to project the multi-head attention scores to the final values.
@@ 2.3. State Space Models
State Space Models (SSMs) are a traditional mathematical framework utilized to depict the dynamic behavior of a system over time (Kalman, 1960). Recent years have found the widespread applications of SSMs in diverse fields like control theory, robotics, and economics (Gu et al., 2021b, a). At its core, SSMs embody the system’s behavior through a collection of hidden variables referred to as ”states”, enabling it to capture temporal data dependencies effectively. Different from RNNs, SSMs are linear models characterised by their associative property. To be specific, in a classical state space model, two key equations are formulated, i.e., state equation and observation equation, to model the relationships between input $x(t)\in\mathbb{R}$ and output $y(t)\in\mathbb{R}$ at current time $t$, through a N-dimensional hidden state $h(t)\in\mathbb{R}^{N}$. The process can be written by
@@ 2.3.1. Discretization
To adhere to the requirements of machine learning settings, SSMs must undergo a process of discretization that transforms continuous parameters into discrete parameters. Discretization methods generally aim to partition continuous time into $K$ discrete intervals with equal integration area as possible. To achieve the goal, as one of the most representative solutions, Zero-Order Hold (ZOH) (Zhang and Chong, 2007; Pechlivanidou and Karampetakis, 2022) is successfully employed in SSMs, which assumes that the function value remains constant over the interval $\Delta=[t_{k-1},t_{k}]$. After ZOH discretization, the SSM equations can be rewritten by
@@ 2.3.2. Convolutional Computation
The discrete SSM, being a linear system, possesses the associated property and therefore integrates seamlessly with convolutional computation. More specifically, it can calculate the output at each time step independently as follows:
@@
where $\mathbf{x}=[x_{0},x_{1},...]$ and $\mathbf{y}=[y_{0},y_{1},...]\in\mathbb{R}^{L}$ denote the input and output sequences, respectively, while $L$ is the sequence length. This convolutional computation allows SSMs to take full advantage of modern matrix computation hardware (e.g., GPUs) to enable parallel computing during training process, which is impossible with RNNs utilizing nonlinear activation functions. Notably, given an input $x(k)$ with $D$ dimensions, the SSM computation will be calculated separately for each dimension to produce a $D$-dimensional output $y(t)$. In this case, the input matrix $\mathbf{B}\in\mathbb{R}^{N\times D}$, the output matrix $\mathbf{C}\in\mathbb{R}^{D\times N}$, and the command matrix $\mathbf{D}\in\mathbb{R}^{D\times D}$, while the state transition matrix remains unchanged, i.e., $\mathbf{A}\in\mathbb{R}^{N\times N}$.
@@ 2.3.3. Relationship among RNN, Transformer, and SSM
The computation algorithms of RNN, Transformer, and SSM are depicted in Figure 2. On the one hand, the conventional RNN operates within a non-linear recurrent framework where each computation depends solely on the previous hidden state and the current input. While this format allows RNNs to quickly generate outputs during auto-regressive inference, it hampers their ability to fully exploit GPU parallel computing, leading to slower model training. On the other hand, the Transformer architecture performs matrix multiplications in parallel across multiple query-key pairs, which can be efficiently distributed across hardware resources, which enables faster training of attention-based models. However, when it comes to generating responses or predictions from Transformer-based models, the inference process can be time-consuming. For instance, the auto-regressive design of language models entails generating each token in the output sequence sequentially, which requires repetitive calculations of attention scores at each step, leading to slower inference times. Unlike RNNs and Transformers, which are limited to supporting only one type of computation, discrete SSMs have the flexibility to support both recurrent and convolutional computations, given its linear property. This unique capability allows SSMs to achieve not only efficient inference but also parallel training. However, it should be noted that the most conventional SSMs are time-invariant, meaning that their $\mathbf{A}$, $\mathbf{B}$, $\mathbf{C}$, and $\Delta$ are unrelated to the model input $x$. This would limit context-aware modeling, which leads to inferior performance of SSMs in certain tasks such as selective copying (Gu and Dao, 2023).

### v3

@@ Abstract.
Deep learning (DL), as a vital technique, has sparked a notable revolution in artificial intelligence (AI), resulting in a great change in human lifestyles. As one of the most representative DL techniques, Transformer architecture has empowered numerous advanced models, especially the large language models (LLMs) that comprise billions of parameters, becoming a cornerstone in deep learning. Despite the impressive achievements, Transformers still face inherent limitations, particularly the time-consuming inference resulting from the quadratic computation complexity of attention calculation. Recently, a novel architecture named Mamba, drawing inspiration from classical state space models (SSMs), has emerged as a promising alternative for building foundation models, delivering comparable modeling abilities to Transformers while preserving near-linear scalability concerning sequence length. This has sparked an increasing number of studies actively exploring Mamba’s potential to achieve impressive performance across diverse domains. Given such rapid evolution, there is a critical need for a systematic review that consolidates existing Mamba-empowered models, offering a comprehensive understanding of this emerging model architecture. In this survey, we therefore conduct an in-depth investigation of recent Mamba-associated studies, covering three main aspects: the advancements of Mamba-based models, the techniques of adapting Mamba to diverse data, and the applications where Mamba can excel. Specifically, we first review the foundational knowledge of various representative deep learning models and the details of Mamba-1&2 as preliminaries. Then, to showcase the significance of Mamba for AI, we comprehensively review the related studies focusing on Mamba models’ architecture design, data adaptability, and applications. Finally, we present a discussion of current limitations and explore various promising research directions to provide deeper insights for future investigations.
State Space Models, Mamba, Sequence Modeling, Foundation Models, Language Models
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2024; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Neural Networks
@@ 1. Introduction
Over the past two decades, deep learning (DL), as the most prominent artificial intelligence (AI) technique, has brought about a revolution in various domains such as healthcare (Jones et al., 2024), autonomous systems (Guan et al., 2024; Fan et al., 2024a), recommender systems (Li et al., 2024a; Zhao et al., 2024b), and financial services (Prata et al., 2024; Zhang et al., 2024d). This period has witnessed the emergence of numerous deep neural networks (DNNs) that have significantly altered human lifestyles, offering immense convenience to individuals. One notable example is U-Net (Ronneberger et al., 2015; Si et al., 2024), a robust deep learning model within the field of vision, which is extensively employed in medical imaging for the examination of radiology scans like MRI and CT scans. Its application assists in the identification and diagnosis of diseases, showcasing its effectiveness in this critical healthcare domain (Williams et al., 2024; Lin et al., 2022). Moreover, Graph Neural Networks (GNNs) are employed in handling graph-structured data to support smart services, such as recommender systems that suggest personalized content, products, or services to users (Fan et al., 2020, 2019b; Wu et al., 2019). Furthermore, Recurrent Neural Networks (RNNs) are extensively adopted in machine translation due to their ability to capture the sequential and contextual information essential for accurate translations (Liu et al., 2014; Su et al., 2017), empowering individuals from diverse linguistic backgrounds to effectively communicate and comprehend each other’s ideas, opinions, and information.
@@
Among the various DL architectures, Transformers have recently stood out and established their dominance across a broad spectrum of applications (Dong et al., 2023; Vert, 2023). For instance, as the most representative large foundation models, large language models (LLMs) like ChatGPT and GPT4 are fundamentally built on the Transformer architecture (Achiam et al., 2023; Qu et al., 2024a; Zhao et al., 2024b). By scaling their model sizes to billions and training on a mix of diverse data sources, these Transformer-based models have demonstrated human-level intelligence with their impressive capabilities in language understanding, common sense reasoning, and in-content-learning (Zhang et al., 2023; Fan et al., 2024b). This remarkable success is bolstered by the attention mechanism (Vaswani et al., 2017), which enables the Transformer-based models to concentrate on relevant parts of the input sequence and facilitate better contextual understanding. However, the attention mechanism also introduces a significant computational overhead that increases quadratically with the input size (Lu et al., 2021; Zhu et al., 2021), which presents challenges in processing lengthy inputs. For example, the rapid growth in computational cost makes Transformers impractical or infeasible to process substantial sequences, thereby limiting their applicability in tasks like document-level machine translation (Maruf et al., 2021) or long document summarization (Koh et al., 2022).
@@
Recently, a promising architecture, structured state space sequence models (SSMs) (Gu et al., 2021a), have emerged to efficiently capture complex dependencies in sequential data, becoming a formidable competitor to Transformer. These models, inspired by classical state space models (Kalman, 1960), can be considered a fusion of recurrent neural networks and convolutional neural networks. They can be computed efficiently using either recurrence or convolution operations, achieving linear or near-linear scaling with sequence length, thus significantly reducing the computational costs. More specifically, as one of the most successful SSM variants, Mamba achieves comparable modeling capabilities to Transformers while maintaining linear scalability with sequence length (Gu and Dao, 2023), propelling it into the realm of focal topics. Mamba first introduces a simple yet effective selection mechanism that enables the model to filter out irrelevant information while retaining necessary and relevant data indefinitely by parameterizing the SSM parameters based on the input. Then, Mamba proposes a hardware-aware algorithm to compute the model recurrently with a scan instead of convolution, achieving up to 3$\times$×faster computation on A100 GPUs. As shown in Figure 1, the powerful modeling capabilities for complex and lengthy sequential data, along with near-linear scalability, position Mamba as an emerging foundation model, poised to revolutionize various domains of research and applications, such as computer vision (Xu et al., 2024b; Zhu et al., 2024), natural language processing (Lieber et al., 2024; Zhao et al., 2024c), healthcare (Ruan and Xiang, 2024; Xing et al., 2024; Wang et al., 2024g), etc. For example, Zhu et al. (2024) propose Vim, which is 2.8$\times$×faster than DeiT (Touvron et al., 2021) and saves 86.8% GPU memory when extracting features for high-resolution images. Dao and Gu (2024) show the connections between SSMs and variants of attention and propose a new architecture that refines selective SSM, achieving 2-8$\times$ faster on language modeling.
@@ 2. Preliminary
Mamba is deeply intertwined with the recurrent framework of Recurrent Neural Networks (RNNs), the parallel computation and attention mechanism of Transformers, and the linear property of State Space Models (SSMs). Therefore, this section aims to present an overview of these three prominent architectures.
2.1. Recurrent Neural Networks (RNNs)
RNNs excel in processing sequential data due to their capability to retain internal memory (Graves and Graves, 2012). Such networks have demonstrated remarkable effectiveness in a multitude of tasks that involve analyzing and predicting sequences, e.g., speech recognition, machine translation, natural language processing, and time-series analysis (Sutskever et al., 2011; Hermans and Schrauwen, 2013). In order to grasp the foundations of recurrent models, this section will offer a brief overview of the standard RNN formulation.
@@ 2.1. Recurrent Neural Networks (RNNs)
However, there are several limitations associated with RNNs. First, RNNs have a restricted capability to effectively extract long-range dynamics within input sequences. As information traverses through successive time steps, the repeated multiplication of weights in the network can lead to dilution or loss of information. Consequently, it becomes challenging for RNNs to retain and recall information from earlier time steps while making predictions. Second, RNNs process sequential data incrementally, restricting their computational efficiency since each time step relies on the preceding one. This makes parallel computations challenging for them. Furthermore, conventional RNNs lack built-in attention mechanisms, which allow the network to capture global information within input sequences. This absence of attention mechanisms hinders the network’s ability to selectively model the crucial segments of the data. To overcome these constraints, Transformers and State Space Models have emerged, each tackling these challenges from different perspectives. These two approaches will be further elaborated upon in the subsequent subsections.
@@ 2.2. Transformers
The Transformer (Vaswani et al., 2017) is a groundbreaking model in the realm of deep learning, revolutionizing various AI applications. Its introduction marked a significant departure from traditional sequence-to-sequence models by employing a self-attention mechanism, facilitating the capture of global dependencies within model inputs. For instance, in natural language processing, this self-attention capability allows the model to comprehend relationships between various positions in a sequence. It achieves this by assigning weights to each position based on its significance relative to other positions. More specifically, a sequence of input vectors x is first transformed into three types of vectors: Query $Q$, Key $K$, and Value $V$ by utilizing linear transformations of the original input, defined by:
@@
Apart from performing a single attention function, multi-head attention is introduced to enhance the model’s ability to capture different types of relationships and provide multiple perspectives on the input sequence. In multi-head attention, an input sequence is processed in parallel by multiple sets of self-attention modules. Each head operates independently, performing the exact computations as in the standard self-attention mechanism. The attention weights from each head are then combined to create a weighted sum of the value vectors. This aggregation step allows the model to leverage information from multiple heads and capture diverse patterns and relationships in the input sequence. Mathematically, the multi-head attention is computed as follows:
(5) $\begin{split}\text{MultiHead}&(Q,K,V)=(S_{1}\oplus S_{2}\oplus...\oplus S_{m})\cdot\mathbf{W}^{O},\\ \text{where}~{}~{}~{}~{}&S_{i}=\text{Softmax}(\frac{Q_{i}K_{i}^{T}}{\sqrt{d_{K}}})V_{i},~{}~{}i\in[1,m],\end{split}$
where $m$ is the number of attention heads, $\oplus$ is the concatenation operation, and $\mathbf{W}^{O}$ is the linear transformation to project the multi-head attention scores to the final values.
@@ 2.3. State Space Models
State Space Models (SSMs) are a traditional mathematical framework utilized to depict the dynamic behavior of a system over time (Kalman, 1960). Recent years have found the widespread applications of SSMs in diverse fields like control theory, robotics, and economics (Gu et al., 2021b, a). At its core, SSMs embody the system’s behavior through a collection of hidden variables referred to as ”states”, enabling it to capture temporal data dependencies effectively. Different from RNNs, SSMs are linear models characterized by their associative properties. To be specific, in a classical state space model, two fundamental equations are formulated, i.e., state equation and observation equation, to model the relationships between input $x(t)\in\mathbb{R}$ and output $y(t)\in\mathbb{R}$ at current time $t$, through a N-dimensional hidden state $h(t)\in\mathbb{R}^{N}$. The process can be written by
@@ 2.3.1. Discretization
To adhere to the requirements of machine learning settings for various real-world scenarios, SSMs must undergo a process of discretization that transforms continuous parameters into discrete parameters. Discretization methods generally aim to partition continuous time into $K$ discrete intervals with equal integration area as possible. To achieve the goal, as one of the most representative solutions, Zero-Order Hold (ZOH) (Zhang and Chong, 2007; Pechlivanidou and Karampetakis, 2022) is successfully employed in SSMs, which assumes that the function value remains constant over the interval $\Delta=[t_{k-1},t_{k}]$. After ZOH discretization, the SSM equations can be rewritten by
@@ 2.3.2. Convolutional Computation
The discrete SSM, being a linear system, possesses the associated property and, therefore, integrates seamlessly with convolutional computation. More specifically, it can calculate the output at each time step independently as follows:
@@
where $\mathbf{x}=[x_{0},x_{1},...]$ and $\mathbf{y}=[y_{0},y_{1},...]\in\mathbb{R}^{L}$ denote the input and output sequences, respectively, while $L$ is the sequence length. This convolutional computation allows SSMs to take full advantage of modern matrix computation hardware (e.g., GPUs) to enable parallel computing during the training process, which is impossible with RNNs utilizing nonlinear activation functions. Notably, given an input $x(k)$ with $D$ dimensions, the SSM computation will be calculated separately for each dimension to produce a $D$-dimensional output $y(t)$. In this case, the input matrix $\mathbf{B}\in\mathbb{R}^{N\times D}$, the output matrix $\mathbf{C}\in\mathbb{R}^{D\times N}$, and the command matrix $\mathbf{D}\in\mathbb{R}^{D\times D}$, while the state transition matrix remains unchanged, i.e., $\mathbf{A}\in\mathbb{R}^{N\times N}$.
@@ 2.3.3. Relationship among RNN, Transformer, and SSM
The computation algorithms of RNN, Transformer, and SSM are depicted in Figure 2. On the one hand, the conventional RNN operates within a non-linear recurrent framework where each computation depends solely on the previous hidden state and the current input. While this format allows RNNs to quickly generate outputs during auto-regressive inference, it hampers their ability to fully exploit GPU parallel computing, leading to slower model training. On the other hand, the Transformer architecture performs matrix multiplications in parallel across multiple query-key pairs, which can be efficiently distributed across hardware resources, which enables faster training of attention-based models. However, when it comes to generating responses or predictions from Transformer-based models, the inference process can be time-consuming. For instance, the auto-regressive design of language models entails generating each token in the output sequence sequentially, which requires repetitive calculations of attention scores at each step, leading to slower inference times. As shown in Table 1, unlike RNNs and Transformers, which are limited to supporting only one type of computation, discrete SSMs have the flexibility to support both recurrent and convolutional computations, given their linear properties. This unique capability allows SSMs to achieve not only efficient inference but also parallel training. However, it should be noted that the most conventional SSMs are time-invariant, meaning that their $\mathbf{A}$, $\mathbf{B}$, $\mathbf{C}$, and $\Delta$ are unrelated to the model input $x$. This would limit context-aware modeling, which leads to inferior performance of SSMs in certain tasks such as selective copying (Gu and Dao, 2023).
Table 1. Pros and cons of three primary architectures-RNNs, Transformers, and SSMs-in auto-regressive sequential modeling tasks.
Comparison RNNs Transformers SSMs
Training Speed Slow (Recurrent) Fast (Parallel) Fast (Convolutional)
Inference Speed Fast (Recurrent) Slow (Quadratic-Time) Fast (Recurrent)
Complexity $O(LD^{2})$ $O(L^{2}D)$ $O(LD^{2})$
Modeling Capabilities   (Hidden State)   (Attention)   (Time-Invariance)